D | Question 18 2 pts The balance of Barkley Cos supplies account was $1,200 at the start of the accounting period. During the year Barkley made two purchases of supplies for $1,400 and $2,200. A physical count of the supplies on hand at the end of the accounting period showed that Barkley had $800 worth of supplies still on hand. For supplies, Barkleys financial statements for this period should include? Supplies of $3,600 on the Balance Sheet Supplies Expense of $800 on the Income Statement Supplies of $4.800 on the Balance Sheet Supplies Expense of $4,000 on the Income Statement
Answer #1

Beginning balance of supplies = 1200

Supplies purchase = 1400+2200 = 3600

Ending balance of supplies = 800

Supplies expense = 1200+3600-800 = 4000

So answer is d) Supplies expense of $4000 on the income statement